Solve the equation in exact differentials:

(y-3x2)dx+(x-4y)dy=0


1
Expert's answer
2022-01-09T10:35:29-0500

We shall solve the equation

"(y-3x^{2})dx +(x-4y)dy"

In exact differentials.

First of all, we shall show that the differential equation is indeed EXACT.

From the equation, we have

"M(x,y)=y-3x^{2}" and

"N(x,y)=x-4y"

We know that for exactness,

"\frac{\partial M(x,y)}{\partial y}" must be equal to "\frac{\partial N(x,y)}{\partial x}"


Now,

"\frac{\partial M(x,y)}{\partial y}=1" and "\frac{\partial N(x,y)}{\partial x}=1"

Showing that the differential equation is EXACT.


We shall then solve the equation,

"F_{x}(x,y)=y-3x^{2}------(1)\\F_{y}(x,y)=x-4y------(2)"

Integrating equ(1) "w.r.t." "x"

"F(x,y)=xy-x^{3}+\phi(y)---(*)"

Differentiating "equ(*)" "w.r.t." "y"

"F_{y}(x,y)=x+\phi^{\prime}(y)----(**)"

Equating "equ(2)" and "equ(**)"

"x-4y=x+\phi^{\prime}(y)\\\implies\phi^{\prime}(y)=-4y\\"

Integrating both sides w.r.t "y"

"\phi(y)=-2y^{2}"


Now, from the equation

"F(x,y)=xy-x^{3}+\phi(y)"


Substituting the value of "\phi(y)=-2y^{2}"


We have

"F(x,y)=xy-x^{3}-2y^{2}"

Which is our solution to the differential equation.